if the supply and the demand of a certain product is given by
D(t)=95-5p(t)+2p'(t) and
S(t)=35-p(t)+3p'(t) respectively where p(t) is the price at any time t. Find the equilibrium price and the long range equilibrium price given that p(0)=50
1
Expert's answer
2021-08-09T15:41:43-0400
"D(t)=S(t)"

"95-5p(t)+2p'(t)=35-p(t)+3p'(t)"

"p'(t) =-4p(t)+60"

"\dfrac{dp}{15-p}=4dt"

Integrate


"\int\dfrac{dp}{15-p}=\int 4dt"


"\ln|15-p|=-4t+\ln C"

"|15-p|=Ce^{-4t}"

"p(0)=50, |15-50|=C"

"C=35"

The equilibrium price 

"p(t)=15+35e^{-4t}"

The long range equilibrium price 


"p=15"